A random sample of n=25 individuals is selected from a population with mean =20,and a treatment is administered to each individual in the sample. After treatment, the sample mean is found to be M =22.2 with as=384. How much difference is there between the mean for the treated sample and the mean for the original population?
Answer #1

25, u=20 ñ = 92.2 i mean difference is X-4 = 22.2-20 = 12.2 Standard error of Sample mean is as follow, UH) = 5² = [ (xi-x)²Critical value Q=0,05 a2= 0.025 ta/2, df = too15,24 =:19064 (osing z-table] at tably rejection /hom I sejection regim rejecti